OpenStudy (anonymous):

Find the value of the discriminant 3x2 - 5x + 6 = 0

OpenStudy (anonymous):

it is in standard form already.

OpenStudy (anonymous):

b^2-4 a c

OpenStudy (anonymous):

well my question says find the discriminant and describe the number and type of roots

OpenStudy (anonymous):

5^2-(4 * 3 * 6)

OpenStudy (anonymous):

-47 so it has 2 imaginary roots

OpenStudy (anonymous):

ohk thnks :)

OpenStudy (anonymous):

are u sure its not 2 complex roots?

OpenStudy (anonymous):

complex roots yes same thing i mean

OpenStudy (anonymous):

wat about x2 + 5x + 6 = 0

OpenStudy (anonymous):

it is standard form also so b^2-4 a c 5^2-(4*6)

OpenStudy (anonymous):

which is 1 so it has 2 real roots :)

OpenStudy (anonymous):

thnk u sooooo very much!!!!!!!!

OpenStudy (anonymous):

i actually have 3 more if u dnt mind helping

OpenStudy (anonymous):

thats if u dnt mind

OpenStudy (anonymous):

remember the rulse is use b^2-4 a c if result=r r for result r>0 it has 2 real roots if r=0 then it has 1 root if r<0 then it has 2 imiginary aka complex roots.

OpenStudy (anonymous):

x2 + 5x + 6 = 0

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

ohk :(

OpenStudy (anonymous):

one more here are your formulas ax^2+bx+c=0 b^2-4 a c youare given x2 + 5x + 6 = 0 a b c so plug that into b^2-4 a c 5^2-4*1*6=1 so it has 2 real roots

OpenStudy (anonymous):

so it has 2 rational roots but wat is the number because it is not 1

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

im dumb at math so excuse me lol but wat does D=?

OpenStudy (anonymous):

D=1 in the very last one

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

im sorry lol i gave u the wrong problem!!! lol i meant to put x2 - 8x +16 = 0... soo sorry!

OpenStudy (anonymous):

i was thinking that :P

OpenStudy (anonymous):

i told u i was sloww lol

OpenStudy (anonymous):

x2 - 8x +16 = 0 a b c b^2-4 a c 8^2-4*16=0 so it has one real root

OpenStudy (anonymous):

and D=0 rite?

OpenStudy (anonymous):

rite!

OpenStudy (anonymous):

thnks hun!